If 56% of American say they live paycheck to paycheck find the probability that five randomly selected American say they live paycheck to paycheck
1
Expert's answer
2020-12-01T06:32:23-0500

The binomial distribution function:

"P(x=k) = nC^k(p^k)(1 – p)^{n-k} \\

n = 5 \\

k = 5 \\

P (x = 5) = 5C^5(0.56^5)(0.44^0) = 0.0551"

The probability that 5 Americans live paycheck to paycheck = 0.0551 or 5.5 %